# HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation

### v2

@@ 3.3 Semantic-aware Hypergraph Encoder Pretraining
Existing HCL methods often overlook node textual semantics and rely on random augmentation strategies, resulting in suboptimal performance (Fang et al. 2024; Wu et al. 2024). Moreover, they struggle to model long-range dependencies, limiting broader contextual understanding (Zhang et al. 2024; Wu et al. 2024). To tackle these issues, we propose a semantic-aware contrastive learning strategy to pre-train $g_{\phi}$. As shown in Figure 1, this stage comprises three components: semantic-aware augmentation, hypergraph encoding, and hierarchical contrastive objectives.
Semantic-aware augmentation. To mitigate the noise from random perturbations, we introduce two novel augmentation techniques: prompt-enhanced text augmentation, which generates diverse yet semantically consistent textual views, and semantic-aware hyperedge drop, which ensures structural perturbations remain aligned with the semantics.
Prompt-enhanced text augmentation. As feature-level perturbations can easily distort semantic meaning, we instead focus on augmenting the raw text using a prompt-enhanced strategy. For each node text $t_{i}$, we construct an augmented view $\tilde{t}_{1,i}=\text{Prompt}(t_{i},\mathcal{H})$, while retaining the original text as $\tilde{t}_{2,i}=t_{i}$. As shown in Figure 2,the prompt template integrates the original text with domain knowledge, global topology, and neighbor context, concatenated into a single sequence. Each component adds a distinct perspective: domain knowledge offers semantic background, global topology introduces structural awareness, and neighbor context provides complementary local semantics. This simple yet effective strategy generates semantically consistent but contextually diverse text variants, with a shared $f_{\theta}$ that maps them to a unified space, denoted as $\tilde{\mathbf{X}}_{1}$ and $\tilde{\mathbf{X}}_{2}$, respectively.
Semantic-aware hyperedge drop. Effective structure augmentation should consider both textual semantics and the original topology. To achieve that, we define a semantic cohesiveness score that reflects both the structural grouping of nodes and their semantic similarity. This score is then used to guide the drop probability of each node–hyperedge incidence: connections within more cohesive hyperedges are more likely to be preserved. In this way, our strategy encourages the retention of structurally and semantically aligned patterns, while selectively pruning noisy or weakly related hyperlinks. We formally define the it as follows.
@@
Given a TAHG $\mathcal{H}(\mathcal{V},\mathcal{E},\mathbf{X})$, the semantic cohesiveness score of a hyperedge $e\in\mathcal{E}$ is defined as the average pairwise cosine similarity among its member nodes. Let $\mathbf{X}_{e}\in\mathbb{R}^{|e|\times d}$ denote the normalized feature matrix for nodes in $e$. The semantic cohesiveness score $s(e)$ is then computed as:
@@
where $\bm{S}_{e}(i,j)$ denotes the similarity between the $i$-th and $j$-th node in hyperedge $e$.
The drop probability of each node-hyperedge incidence $(v_{i},e_{j})$ where $v_{i}\in e_{j}$ is then guided by $s(e_{j})$:
@@
where $\sigma(\cdot)$ is the sigmoid function and $\tau_{\text{drop}}$ is a temperature hyperparameter. Higher cohesiveness scores lead to lower drop probabilities. In implementation, we construct a binary masking matrix $\mathbf{M}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where each entry $m_{ij}\sim\mathcal{B}(1-p_{\text{drop}}(v_{i},e_{j}))$ is sampled from a Bernoulli distribution. The perturbed incidence matrix $\tilde{\mathbf{H}}$ is represented as:
@@
Hypergraph encoder. We take the HGNN (Feng et al. 2019) with the element-wise mean pooling layer as our hypergraph encoder. At the $l$-th layer, the hyperedge representation $\mathbf{Z}_{\mathcal{E}}^{(l)}\in\mathbb{R}^{|\mathcal{E}|\times d}$ and the node representation $\mathbf{Z}_{\mathcal{V}}^{(l)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ are updated as:
@@
where $\rho(\cdot)$ is the PReLU activation function, $\mathbf{\Theta}_{\mathcal{E}}^{(l)}$ and $\mathbf{\Theta}_{\mathcal{V}}^{(l)}$ are trainable weight matrices, and $\bm{b}_{e}^{(l)}$ and $\bm{b}_{v}^{(l)}$ are corresponding bias terms. $\mathbf{W}$ is initialized as an identity matrix, which means equal weights for all hyperedges.
Hierarchical contrastive objectives. Prior works (Lee and Shin 2023; Li et al. 2024) have demonstrated that multi-scale contrastive objectives are beneficial for HSSL. To better capture structural semantics at different granularities, we adopt a hierarchical contrastive learning objective spanning node, hyperedge, and subgraph levels. In particular, the novel subgraph-level objective, based on $s$-walk sampling, enables the model to capture long-range dependencies beyond local structure. All objectives are optimized with InfoNCE loss (Oord, Li, and Vinyals 2018).
@@
where $\tau_{n}$ is the temperature coefficient. In practice, we symmetrize the loss by using the second view’s embedding as the anchor and obtain the final node-level contrastive loss:
@@
where $\tau_{e}$ is the temperature coefficient. Symmetrizing the loss, the final hyperedge-level objective can be written as:
@@
Figure 3: An example of the $s$-walk-based subgraph sampling process. For center node $v_{c}$, we perform $s$-walk with $s=2$ and $l=4$ on $\mathcal{H}$. The hyperedge sequences visited are $e_{1},e_{3},e_{4},e_{5}]$. The sampled subgraph is denoted as $\mathcal{H}_{s_{c}}$.
While node- and hyperedge-level objectives capture local semantic and structural signals, they may fail to model long-range dependencies in hypergraphs. To bridge this gap, we further introduce a subgraph-level objective by constructing broader structural signals through a novel $s$-walk-based sampling strategy. Unlike random walk, $s$-walk is a high-order sequence in hypergraphs, where $s$ measures the overlap strength between hyperedges, enabling the extraction of cohesive and semantically meaningful substructures unique to hypergraphs (Aksoy et al. 2020; Preti, De Francisci Morales, and Bonchi 2024). We next define $s$-walk and detail our subgraph sampling method.
@@
Given $s\in\mathbb{N}^{+}$, an $s$-walk of length $l$ is a sequence of hyperedges $e_{1},\ldots,e_{l}]$ such that each consecutive pair of hyperedges shares at least $s$ nodes, i.e., $|e_{j}\cap e_{j+1}|\geq s$ for all $j\in[1,l-1]$.
As illustrated in Figure 3, we construct an $s$-walk-based subgraph $\mathcal{H}_{s_{c}}=\{\mathcal{V}_{s_{c}},\mathcal{E}_{s_{c}}\}$ by first selecting a starting hyperedge $e_{1}$ that contains the center node $v_{c}$, and then performing an $s$-walk of length $l$. We randomly sample one node from each visited hyperedge. The sampled subgraph includes all visited hyperedges and their associated nodes: $\mathcal{E}_{s_{c}}=\{e_{j}\}_{j=1}^{l}$, $\mathcal{V}_{s_{c}}=\bigcup_{j=1}^{l}e_{j}$. Each subgraph $\mathcal{H}_{s_{c}}$ can be encoded by the shared hypergraph encoder $g_{\phi}(\cdot)$, and its representation is obtained via mean pooling $P_{\text{mean}}$ over the node embeddings:
@@
However, constructing and encoding subgraphs for all nodes incurs a high computational cost. To mitigate this, we follow prior findings suggesting that high-degree nodes tend to carry richer contextual information (Fang et al. 2025). Specifically, we select a subset of representative nodes $\mathcal{V}_{r}\subseteq\mathcal{V}$ by ranking node degrees and retaining the top $r\%$. As validated in our experiments, this strategy maintains competitive performance even with a significantly reduced number of subgraphs. For each anchor node $v_{m}\in\mathcal{V}_{r}$, we obtain subgraph representations $\mathbf{z}_{1,s_{m}}$ and $\mathbf{z}_{2,s_{m}}$ from two augmented views. Similar to the former contrastive signals, the subgraph-level loss with anchor $\mathbf{z}_{1,s_{m}}$ is defined as:
@@
Overall objective. Finally, by integrating Eq. (10), (12), and (15), our proposed contrastive loss is formulated as:
@@
where $\lambda_{e}$ and $\lambda_{s}$ are balancing weights. This hierarchical contrastive framework enables the model to capture patterns at multiple granularities, generating more robust and discriminative node representations.
@@ 4 Experiments
Graph SSL
GraphCL + SE 59.63 ± 1.58 56.69 ± 1.99 62.86 ± 0.22 42.61 ± 0.21 34.45 ± 0.38 52.11 ± 0.43
GraphCL + GIANT 57.58 ± 1.90 55.64 ± 1.50 62.08 ± 0.24 42.74 ± 0.22 34.65 + 0.23 52.04 ± 0.34
@@
VilLain (w/o feat.) 42.22 ± 1.72 53.72 ± 1.00 40.99 ± 3.06 33.94 ± 1.39 20.99 ± 0.91 OOM
HyperBERT 30.25 ± 3.04 26.85 ± 0.66 OOM OOM OOM OOM
HiTeC 66.38 ± 1.73 74.07 ± 0.88 79.81 ± 0.16 59.64 ± 0.16 55.81 ± 0.12 67.18 ± 0.19
Table 2: Node classification accuracy (% mean ± std) across six datasets. The best and second-best results are highlighted in bold and underline, respectively. “w/o feat.” means without node features. OOM denotes out-of-memory on a 24GB GPU.
@@ 4.1 Experimental Setup
Datasets. Since there are no publicly available TAHG datasets, we introduce six new text-attributed hypergraph datasets using a max clique-based hypergraph reconstruction method (Bron and Kerbosch 1973; Wang and Kleinberg 2024). These include two co-citation hypergraphs: Citeseer (Giles, Bollacker, and Lawrence 1998) and Cora (Sen et al. 2008), and four co-purchasing networks: History, Photo, Computers, and Fitness (Yan et al. 2023). Detailed dataset statistics are presented in Table 1, with comprehensive descriptions provided in Appendix A.1.
Baselines. We compare HiTeC with seven strong SSL baselines covering both graph- and hypergraph-based methods. Since these methods cannot handle raw text directly, we extract features using shallow encoders (SE), PLMs (BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019)), and graph-agnositc encoder (GIANT (Chien et al. 2022a)). Full details are provided in Appendix A.2.
Evaluation protocol. We evaluate our approach on two hypergraph benchmark tasks: node classification and hyperedge prediction. After pretraining with HiTeC, we follow the linear evaluation protocol (Lee and Shin 2023; Li et al. 2024; Kim et al. 2024a), where the learned node embeddings are fixed and used as inputs to downstream classifiers. Full evaluation details are provided in Appendix A.3.
Implementation details. All baselines use official implementations with default settings. For HiTeC, we tune the contrastive loss weights $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$. The overlap size $s$ is selected from $\{1,2,3,4,5,10\}$, and the sampling ratio rr% from $\{10\%,20\%,30\%,40\%,50\%\}$. Full implementation details are provided in Appendix A.4.

### v3

@@ 3.2 Semantic-aware Hypergraph Encoder Pretraining
Existing HCL methods typically overlook node textual semantics and rely on random augmentation strategies, resulting in suboptimal performance Fang et al. (2024); Wu et al. (2024). Moreover, they struggle to model long-range dependencies, limiting broader contextual understanding Zhang et al. (2024); Wu et al. (2024). To tackle these issues, we propose a semantic-aware contrastive learning strategy to pre-train $g_{\phi}$. As shown in Figure 1, this stage comprises three components: semantic-aware augmentation, hypergraph encoding, and hierarchical contrastive objectives.
Semantic-aware augmentation. To eliminate the impact of noise arising from random perturbation, we introduce two semantic-aware augmentation strategies: structure-contextualized text augmentation and semantic-aware hyperedge dropping.
Structure-contextualized text augmentation. As perturbations applied directly to text embeddings can easily distort semantic meaning, we instead augment raw text with structure-aware contextual information. Specifically, for each node text attribute $t_{i}$, we first construct a structure-contextualized text view by concatenating three contextual components with the original text: $\tilde{t}_{1,i}=[\texttt{[DK]}\,||\,\texttt{[GT]}\,||\,\texttt{[NC]}\,||\,t_{i}]$, where: (i) [DK] denotes domain knowledge (e.g., academic or e-commerce) shared across all nodes; (ii) [GT] encodes global topological information for node $v_{i}$ (e.g., node degree and local sparsity); and (iii) [NC] represents neighbor context, constructed by concatenating the raw texts of sampled neighbors of $v_{i}$. The second text view is the original text, i.e., $\tilde{t}_{2,i}=t_{i}$. This simple yet effective strategy generates semantically consistent but contextually diverse text variants, which are encoded by the shared encoder $f_{\theta}$ into a unified embedding space, denoted as $\tilde{\mathbf{X}}_{1}$ and $\tilde{\mathbf{X}}_{2}$, respectively. The full template for constructing the structure-contextualized text view is provided in Appendix A.
Semantic-aware hyperedge dropping. Effective structure augmentation should consider both textual semantics and topology. To achieve this, we define a semantic cohesiveness score that reflects both the structural grouping of nodes and their semantic similarity, which is further used to guide the drop probability of each node–hyperedge incidence. Connections within more cohesive hyperedges are more likely to be preserved, encouraging the retention of structurally and semantically aligned patterns, while selectively pruning noisy or weakly node–hyperedge connections.
@@
Given a TAHG $\mathcal{H}(\mathcal{V},\mathcal{E},\mathbf{X})$, the semantic cohesiveness score of a hyperedge $e\in\mathcal{E}$ is defined as the average pairwise similarity among its member nodes. Let $\mathbf{X}_{e}\in\mathbb{R}^{|e|\times d}$ denote the normalized feature matrix for nodes in $e$. The semantic cohesiveness score $s(e)$ is then computed as:
@@
where $\bm{S}_{e}(i,j)$ denotes the similarity between the $i$-th and $j$-th nodes in hyperedge $e$.
The drop probability of each incidence $(v_{i},e_{j})$, where $v_{i}\in e_{j}$, is then determined by $s(e_{j})$:
@@
where $\sigma(\cdot)$ is the sigmoid function and $\tau_{\text{drop}}$ is a temperature hyperparameter. Higher cohesiveness scores correspond to lower drop probabilities. In implementation, we construct a binary masking matrix $\mathbf{M}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where entries $m_{ij}\sim\mathcal{B}(1-p_{\text{drop}}(v_{i},e_{j}))$ are sampled from a Bernoulli distribution. The perturbed matrix $\tilde{\mathbf{H}}$ is:
@@
Hypergraph encoder. We take the HGNN Feng et al. (2019b) with the element-wise mean pooling layer as our hypergraph encoder. At the $l$-th layer, the hyperedge representation $\mathbf{Z}_{\mathcal{E}}^{(l)}\in\mathbb{R}^{|\mathcal{E}|\times d}$ and the node representation $\mathbf{Z}_{\mathcal{V}}^{(l)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ are updated as:
@@
where $\rho(\cdot)$ is the PReLU activation function, $\mathbf{\Theta}_{\mathcal{E}}^{(l)}$ and $\mathbf{\Theta}_{\mathcal{V}}^{(l)}$ are trainable weight matrices, and $\bm{b}_{e}^{(l)}$ and $\bm{b}_{v}^{(l)}$ are bias terms. $\mathbf{W}$ is initialized as an identity matrix, assigning equal weights to all hyperedges.
Hierarchical contrastive objectives. Prior works Lee and Shin (2023); Li et al. (2024) have shown that multi-scale contrastive objectives are beneficial for HSSL. To better capture structural semantics at different granularities, we adopt a hierarchical contrastive learning framework across node, hyperedge, and subgraph levels. In particular, we introduce a novel subgraph-level objective based on $s$-walk sampling to capture long-range dependencies beyond local structures. All objectives are optimized using the widely adopted contrastive loss, InfoNCE Oord et al. (2018).
@@
where $\tau_{n}$ is the temperature coefficient. In practice, we further symmetrize the contrastive loss by using the second view’s embedding as the anchor, yielding the final node-level contrastive objective:
@@
where $\tau_{e}$ is the temperature coefficient. By symmetrizing the contrastive loss, the final hyperedge-level objective can be written as:
@@
Figure 2: An example of the $s$-walk-based subgraph sampling process. Given the center node $v_{c}$, we perform an $s$-walk with $s=2$ and $l=3$ on $\mathcal{H}$. The resulting hyperedge traversal sequence is $e_{1},e_{3},e_{4}]$, and the sampled subgraph is denoted as $\mathcal{H}_{s_{c}}$.
While node- and hyperedge-level objectives effectively capture local structural signals, they remain insufficient for modeling long-range dependencies in hypergraphs. We therefore introduce a subgraph-level objective based on an $s$-walk sampling strategy. Unlike random walks, an $s$-walk defines a high-order traversal sequence in hypergraphs, where $s$ controls the overlap strength between consecutive hyperedges, enabling the extraction of cohesive and semantically meaningful substructures unique to hypergraphs Aksoy et al. (2020); Preti et al. (2024). We next formally define $s$-walk and describe our subgraph sampling method in detail.
@@
Given $s\in\mathbb{N}^{+}$, an $s$-walk of length $l$ is defined as an ordered sequence of hyperedges $e_{1},\ldots,e_{l}]$, where every consecutive pair of hyperedges shares at least $s$ common nodes, i.e., $|e_{j}\cap e_{j+1}|\geq s$ for all $j\in[1,l-1]$.
As illustrated in Figure 2, an $s$-walk-based subgraph $\mathcal{H}_{s_{c}}=(\mathcal{V}_{s_{c}},\mathcal{E}_{s_{c}})$ is constructed from a center node $v_{c}$ via iterative $s$-walk sampling. We first randomly select a hyperedge $e_{1}$ containing $v_{c}$. At each step, we identify hyperedges sharing at least $s$ nodes with the current hyperedge $e_{j}$, randomly sample one as $e_{j+1}$, and then randomly choose a node from $e_{j+1}$ as the transition node. The process stops when the traversal length reaches $l$. The sampled subgraph consists of the visited hyperedges $\mathcal{E}_{s_{c}}=\{e_{j}\}_{j=1}^{l}$ and their associated nodes $\mathcal{V}_{s_{c}}=\bigcup_{j=1}^{l}e_{j}$. Each sampled subgraph $\mathcal{H}_{s_{c}}$ is encoded by the shared hypergraph encoder $g_{\phi}(\cdot)$, and its representation is obtained through mean pooling $P_{\text{mean}}$ over the corresponding node embeddings:
@@
However, constructing and encoding subgraphs for all nodes incurs a high computational cost. To mitigate this, we follow prior findings suggesting that high-degree nodes tend to carry richer contextual information Fang et al. (2025). Specifically, we select a representative node subset $\mathcal{V}_{r}\subseteq\mathcal{V}$ by ranking node degrees and retaining the top $r\%$. As validated in our experiments, this strategy maintains competitive performance even with a significantly reduced number of subgraphs. For each anchor node $v_{m}\in\mathcal{V}_{r}$, we obtain subgraph representations $\mathbf{z}_{1,s_{m}}$ and $\mathbf{z}_{2,s_{m}}$ from two augmented views. Similar to the former contrastive signals, the subgraph-level loss with anchor $\mathbf{z}_{1,s_{m}}$ is defined as:
@@
Overall objective. Combining Eq. (10), (12), and (15), the overall contrastive loss is defined as: contrastive loss is defined as:
@@
where $\lambda_{e}$ and $\lambda_{s}$ are balancing weights. A detailed complexity analysis is provided in Appendix B.
@@ 4 Experiments
Graph SSL GraphCL + SE 59.63 ± 1.58 56.69 ± 1.99 62.86 ± 0.22 42.61 ± 0.21 34.45 ± 0.38 52.11 ± 0.43
GraphCL + GIANT 57.58 ± 1.90 55.64 ± 1.50 62.08 ± 0.24 42.74 ± 0.22 34.65 ± 0.23 52.04 ± 0.34
@@
VilLain (w/o feat.) 42.22 ± 1.72 53.72 ± 1.00 40.99 ± 3.06 33.94 ± 1.39 20.99 ± 0.91 42.04 ± 0.04
HyperBERT 58.25 ± 3.04 47.03 ± 0.77 62.12 ± 0.19 41.46 ± 0.14 26.79 ± 0.20 OOM
HiTeC 68.69 ± 0.97 74.07 ± 0.88 79.81 ± 0.16 59.64 ± 0.16 55.81 ± 0.12 67.18 ± 0.19
Table 2: Evaluation results on node classification task(% mean $\pm$ std). The best and second-best results are highlighted in bold and underline, respectively. “w/o feat.” denotes methods without node features. OOM indicates that the method runs out of memory on a single 80GB GPU.
@@ 4.1 Experimental Setup
Datasets. As no public TAHG datasets are available, we construct six text-attributed hypergraph datasets via max clique-based hypergraph reconstruction Bron and Kerbosch (1973); Wang and Kleinberg (2024). They include two co-citation hypergraphs (Citeseer Giles et al. (1998) and Cora Sen et al. (2008)) and four co-purchasing networks (History, Photo, Computers, and Fitness Yan et al. (2023)). Dataset statistics are summarized in Table 1, with details in Appendix C.1.
Baselines. We compare HiTeC with eight representative SSL baselines: (1) three graph-based methods (GraphCL You et al. (2020), BGRL Thakoor et al. (2022), and GraphMAE Hou et al. (2022)); (2) four hypergraph-based methods (TriCL Lee and Shin (2023), SE-HSSL Li et al. (2024), HypeBoy Kim et al. (2024a), and VilLain Lee et al. (2024)); and (3) one TAHG-based method, HyperBERT Bazaga et al. (2024). Since graph- and hypergraph-based methods cannot process raw text directly, we extract node features using different text encoders. Specifically, both graph- and hypergraph-based methods use shallow encoders (SE), i.e., Skip-Gram Mikolov et al. (2013). Graph-based methods additionally use GIANT Chien et al. (2022a), a self-supervised encoder for text-attributed graphs (TAGs), while hypergraph-based methods use pretrained language models, including BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019). Detailed descriptions of the baselines are provided in Appendix C.2.
Evaluation protocol. We evaluate our approach on two hypergraph benchmark tasks: node classification and hyperedge prediction. After pretraining with HiTeC, we adopt the linear evaluation protocol Lee and Shin (2023); Li et al. (2024); Kim et al. (2024a), where the learned node embeddings are frozen and used as input features for downstream classifiers. For node classification, we train a logistic regression classifier with a 10%/10%/80% train/validation/test split. For hyperedge prediction, following Kim et al. (2024a), we sample an equal number of negative hyperedges to match the ground-truth positives and split the data using a 60%/20%/20% ratio, after which a two-layer MLP classifier is trained for prediction. For each dataset, we report the average test accuracy and standard deviation over 20 random data splits, each evaluated with 5 different random initializations. Full evaluation details are provided in Appendix C.3.
Implementation details. For HiTeC, we adopt BERT-Base Devlin et al. (2019) as the default text encoder and apply LoRA Hu et al. (2022) for parameter-efficient fine-tuning. We further evaluate other text encoder backbones in Appendix D.2. We tune the $s$-walk length $l$ and overlap size $s$ in $\{1,2,3,4,5,10\}$, the sampling ratio $r$ in $\{10,20,30,40,50\}$, and $\lambda_{e}$ and $\lambda_{s}$ in $\{1,2,3,4\}$. Additional implementation details are shown in Appendix C.4.